# April 2026

### 05 Apr 2026 v1

@@ Abstract
This is the second and final part of a work initiated in [Mor26]. We construct a class of positive-energy relativistic spatial localization observables in Minkowski spacetime within the standard framework of quantum field theory, based on the stress–energy–momentum tensor smeared with suitable test functions. For every fixed timelike direction, the construction yields a family of positive operator-valued measures (POVMs) defined on spacelike hypersurfaces, which are well defined on each $n$-particle sector and satisfy a natural relativistic causality condition ruling out superluminal propagation of detection probabilities. The proposed localization observables arise from local or quasi-local quantum field theoretic quantities, thereby providing a rigorous realization of previously heuristic constructions. In the one-particle sector, the scheme reduces to the observable introduced in [Mor23], and its first moment reproduces the Newton–Wigner position operator under suitable normalization conditions. Since the normally ordered stress–energy–momentum tensor fails to be positive on the full Fock space due to the Reeh–Schlieder theorem, we analyze the role of quantum energy inequalities and establish lower bounds that allow us to control deviations from positivity. This leads to the introduction of regularized families of bounded from below operators approximating the localization effects. We further construct conditional localization observables associated with finite laboratories by means of suitably modified local energy operators and their Friedrichs self-adjoint extensions. In particular, thanks to Haag duality and a result by Kadison on the affiliation of Friedrichs extensions with von Neumann algebras, the resulting conditional POVMs are shown to belong to local von Neumann algebras, and thus to commute when associated with causally separated regions, in agreement with the Araki–Haag–Kastler framework. Our results provide a quantum field theoretic implementation of the idea that commutativity of localization observables should be recovered at the level of conditional measurements in spacetime regions of finite extent.
@@ 1.1 Localization, Causality, Commutativity, and QFT
where $V$ is a given unitary operator, has the properties of a conditional localization POVM for states whose probability of finding the system in $\Delta_{0}$ (measured with the effect $VA(\Delta_{0})V^{\dagger}$) is close to $1$. This intepretation makes sense in the general case where the operators $A(\Delta)$ are positive and define a positive-operator valued measure on $\Sigma$, but are not necessarily effects, i.e., bounded by $I$.
@@
For future convenience, observe that, if ${\mathbb{R}}^{4}\ni x\mapsto U_{x}$ is the unitary representation of the translation group of ${\mathbb{M}}^{4}$, then $U_{x}:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]U_{x}^{\dagger}=:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}]$ where $f_{x}(y):=f(y-x)$. Concerning the generalization developed in this work, we consider an $n$-particle state $\Psi\in{{\cal H}}^{(n)}$ with $n>0$. For every Borel set $\Delta\subset\Sigma$, where $\Sigma$ is a spacelike $3$-plane of Minkowski spacetime $\mathbb{M}$}, the effects $A_{f}^{u}(\Delta)$ of the notion of spatial localization we construct satisfy
@@
The family $A^{u}_{f}(\Delta)$ is a well-behaved positive-energy localization observable in each space ${{\cal H}}^{(n)}$, in particular for one-particle states. Notably, the integral for $\Delta=\Sigma$ is normalized independently of the choice of $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$, provided $\int_{\mathbb{M}}f^{2}d^{4}x=1$. The causality requirement denoted by CC and already established in [Mor23, DRM24] also holds for the localization observables constructed here. We further analyze the physical meaning of the notion of localization introduced above, showing that it is associated with the center of $u$-energy of the field and that, for one-particle states, it reduces to the observable already introduced in [Mor23]. In the one-particle case, one can say even more: if one takes $f(x^{0},\vec{x})=h^{\prime}(x^{0})h(\vec{x})$, then, in the non-relativistic large-mass limit, $h$ determines the precision of a von Neumann measurement scheme for detecting a single particle, whereas $h^{\prime}$ can be chosen independently.
@@ 1.3 Notions and notations
A. Minkowski spacetime. A four-dimensional real affine space whose space of translations $\mathsf{V}$ is equipped with a bilinear, non-degenerate, symmetric form $g$ with signature (,+,+,+)(-,+,+,+) is the Minkowski spacetime $\mathbb{M}$}. The points of $\mathbb{M}$} are called events and $g$ is called the Minkowski metric. We shall make use of the notation $u\cdot v:=g(u,v)$ if $u,v\in\mathsf{V}$. We also use the dot to indicate the standard (positive) scalar product of $3$-vectors $\vec{u}\cdot\vec{v}$, viewing them as spacelike vectors (see below). Upon choosing an origin $o\in{\mathbb{M}}$, the points $p\in{\mathbb{M}}$ are in one-to-one correspondence with vectors of $\mathsf{V}$ through the map ${\mathbb{M}}\ni p\mapsto p-o\in\mathsf{V}$. We shall take advantage of this identification several times in the rest of this paper. If $p\in{\mathbb{M}}$ and $v\in\mathsf{V}$, $q=p+v$ means that qp=vq-p=v.
@@
If $A\subset{\mathbb{M}}$, the causal future $J^{+}(A):=\{p\in{\mathbb{M}}\>|\>p-q\in\overline{\mathsf{V}_{+}}\quad\mbox{for some $q\in A$}\}$ represents the events of $\mathbb{M}$} in the future of $A$ that can be physically influenced by $A$. The causal past $J^{-}(A):=\{p\in{\mathbb{M}}\>|\>q-p\in\overline{\mathsf{V}_{+}}\quad\mbox{for some $q\in A$}\}$ is defined symmetrically.
@@
A rest space of a Minkowski reference frame $u\in{\mathsf{T}}_{+}$ is an affine $3$-plane $\Sigma\subset{\mathbb{M}}$ $g$-normal to $u$, written $u\perp\Sigma$. If $o\in{\mathbb{M}}$ is a given origin, the family of rest spaces of $u\in{\mathsf{T}}_{+}$ is labeled by the time at which they occur in the said reference frame, $t_{o,\Sigma}=-(p-o)\cdot u$, which does not depend on $p\in\Sigma$. Every spacelike affine $3$-plane is the rest space of a Minkowskian reference frame, indicated by $u_{\Sigma}$ and said to be adapted to $\Sigma$, at some time. $u_{\Sigma}$ is the future-directed unit normal vector to $\Sigma$, which is necessarily timelike.
@@
$\mathscr{B}(\Sigma)$ denotes the family of Borel subsets of a spacelike $3$-plane $\Sigma$ and $\mathscr{B}_{b}(\Sigma)\subset\mathscr{B}(\Sigma)$ the subfamily of bounded elements, boundedness being equivalently referred to any of the said coordinate systems. $d\Sigma$ denotes the natural translationally invariant Borel measure on the rest space $\Sigma$ of $u_{\Sigma}$, which coincides with the Lebesgue measure on the ${\mathbb{R}}^{3}$ space of the spatial coordinates of any Minkowskian coordinate system $x^{0},x^{1},x^{2},x^{3}$ comoving with $u_{\Sigma}$. It is easy to prove that $d\Sigma$ does not depend on the choice of such a Minkowskian coordinate system. We use the notation $|\Delta|=\int_{\Sigma}\chi_{\Delta}(x)d\Sigma(x)=\int_{{\mathbb{R}}^{3}}\chi_{\Delta}(x)d^{3}x$ for the Lebesgue measure of $\Delta\in\mathscr{B}(\Sigma)$.
@@
Operators in Hilbert space have their own domains, $A:D(A)\to{{\cal H}}$, where $D(A)\subset{{\cal H}}$ is a subspace. The operations $A+B$, ABAB, aAaA ($a\in{\mathbb{C}}$) are defined on their standard domains: $D(A+B):=D(A)\cap D(B)$, $D(AB):=\{x\in D(B)\>|\>Bx\in D(A)\}$, $D(aA):=D(A)$ unless $a=0$, in which case $D(0A):=D(0):={{\cal H}}$. $A\geq 0$ means that $\langle\psi|A\psi\rangle\geq 0$ if $\psi\in D(A)$. In that case we say that the operator $A$ is positive.
@@ 2.1 Localization observables, causal conditions, non-commutativity
For future convenience, observe that, if $V\in{\mathfrak{B}}({{\cal H}})$ is a unitary operator and $({{\cal H}},{\cal R},A,U)$ is a relativistic spatial localization observable, then $({{\cal H}},{\cal R},A_{V},U_{V})$ is, where $A^{v}_{V}(\Delta):=VA^{v}(\Delta)V^{\dagger}$ and $(U_{V})_{g}:=VU_{g}V^{\dagger}$ is another relativistic spatial localization observable. It also satisfies the causality condition below and is of positive-energy type if $({{\cal H}},{\cal R},A,U)$ does.
@@
A relativistic spatial localization observable $({{\cal H}},{\cal R},A,U)$ is causal if it satisfies the following. For every $\Delta\in{\cal R}$, every spacelike $3$-plane $\Sigma$, and every $v\in{\mathsf{S}}$,
@@
The explicit examples of relativistic spatial localization observables, especially of positive-energy type, presented in [Cas17, Mor23, DRM24, Cas24, Cas25, CDM26] for various types of particles satisfy CC. They therefore show that Hegerfeldt’s causality issue can in fact be regarded as harmless when one considers certain unsharp notions of localization, whereas sharp ones are ruled out.
In [Mor26], after an accurate analysis of relativistic locality from Busch’s perspective, we have asserted that, dealing with relativistic spatial localization observable (also satisfying CC), there is no compelling reason for requiring that $A^{v}(\Delta),A^{v^{\prime}}(\Delta^{\prime})]=0$ when $\Delta$ and $\Delta^{\prime}$ are sharply causally separated, i.e., they are included in respective open regions which are causally separated. This is true for systems like particles which have the propensity to localize in a unique position of a rest space when the rest space is ideally filled with detectors. The found result is in agreement with the celebrated achievement by Halvorson and Clifton [HaCl02] which proves that the above commutativity is actually forbidden if $({{\cal H}},{\cal R},A,U)$ is of positive-energy type [^4] This is even true with a weaker notion of localization observable than the one of Definition 2.2, see the review in [Mor26].. Failure of commutativity has an important consequence. If $\Delta$ and $\Delta^{\prime}$ are bounded regions contained in causally separated open sets $\cal O$} and $\cal O$$^{\prime}$ of $\mathbb{M}$}, one concludes that $A^{v}(\Delta)$ and $A^{v}(\Delta^{\prime})$ cannot belong to the corresponding local algebras of observables in the AHK approach. Indeed, if this were the case, they would necessarily commute. This shows that the effects $A^{v}(\Delta)$ of a relativistic spatial localization observable are not local observables in AHK sense.
@@ 2.2 Conditional localization and commutativity
If $\Sigma$ is a flat $3$-dimensional plane, a laboratory $L(\Delta_{0})\subset{\mathbb{M}}$, with space given by a bounded set $\Delta_{0}\in\mathscr{B}(\Sigma)$, is defined as the causal completion $L(\Delta_{0}):=(\Delta_{0}^{\perp\!\!\!\!\perp})^{\perp\!\!\!\!\perp}$. If $\Delta_{0}$ is open in the relative topology, the resulting open set $L(\Delta_{0})$ is a globally hyperbolic spacetime in its own right, with $\Delta_{0}$ as a possible smooth spacelike Cauchy surface, as well as other curved smooth spacelike Cauchy surfaces.
@@
In other words, if we start with a state $\rho\in{\mathsf{S}}({{\cal H}})$ whose probability of finding the system in $\Delta_{0}$ is close to $1$, $B^{V}_{\Delta_{0}}(\Delta)$ measures the fraction of detections in $\Delta$ over the total number of detections observed in the laboratory $\Delta_{0}$, when an ensemble of identical systems is prepared in the initial state $\rho$} and the relativistic spatial localization observable with effects ${\cal R}\ni\Delta\mapsto A_{V}(\Delta):=VA(\Delta)V^{\dagger}$ is used to define the spatial localization of the system. In this sense, $B^{V}_{\Delta_{0}}$ represents, at least in the above limit, a conditional localization observable: $B^{V}_{\Delta_{0}}(\Delta)$ gives the probability of finding the system in $\Delta\subset\Delta_{0}$, provided that it is known to be found in the bounded measurable set $\Delta_{0}$. This interpretation becomes increasingly accurate as the initial probability of finding the system in $\Delta_{0}$ approaches unity.
@@ 3.1 Bosonic Fock space and particle states
The latter identity is valid in every Minkowskian coordinate frame, where one identifies ${\mathbb{R}}^{3}$ with the $3$-space of the spatial part of the four-momenta, and

### 07 Apr 2026 v2

@@ Abstract
This paper constitutes the second and final part of a work initiated in [Mor26]. We construct a class of positive-energy relativistic spatial localization observables in Minkowski spacetime within the standard framework of quantum field theory, based on the stress–energy–momentum tensor smeared with suitable test functions. For each fixed timelike direction, the construction yields a family of positive operator-valued measures (POVMs) defined on spacelike hypersurfaces, which are well defined on every $n$-particle sector and satisfy a natural relativistic causality condition ruling out superluminal propagation of detection probabilities. The proposed localization observables arise from local or quasi-local quantum field theoretic quantities, thereby providing a rigorous realization of previously heuristic constructions. In the one-particle sector, the scheme reduces to the observable introduced in [Mor23], and its first moment reproduces the Newton–Wigner position operator under suitable normalization conditions. Since the normally ordered stress–energy–momentum tensor need not be positive on the full Fock space, in view of the Reeh–Schlieder theorem, we analyze the role of quantum energy inequalities and establish lower bounds that allow us to control deviations from positivity. This leads to the introduction of regularized families of operators bounded from below that approximate the localization effects. We further construct conditional localization observables associated with finite laboratories by means of suitably modified local energy operators and their Friedrichs self-adjoint extensions. In particular, by Haag duality and a result of Kadison on the affiliation of Friedrichs extensions with von Neumann algebras, the resulting conditional POVMs are shown to belong to local von Neumann algebras and hence to commute when associated with causally separated regions, in agreement with the Araki–Haag–Kastler framework. Our results provide a quantum field theoretic implementation of the idea that commutativity of localization observables should be recovered at the level of conditional measurements in spacetime regions of finite extent.
@@ 1.1 Localization, Causality, Commutativity, and QFT
where $V$ is a given unitary operator, has the properties of a conditional localization POVM for states whose probability of finding the system in $\Delta_{0}$ (measured with the effect $VA(\Delta_{0})V^{\dagger}$) is close to 11. This intepretation makes sense in the general case where the operators $A(\Delta)$ are positive and define a positive-operator valued measure on $\Sigma$, but are not necessarily effects, i.e., bounded by II.
@@
For future convenience, observe that, if ${\mathbb{R}}^{4}\ni x\mapsto U_{x}$ is the unitary representation of the translation group of ${\mathbb{M}}^{4}$, then $U_{x}:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]U_{x}^{\dagger}=:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}]$ where $f_{x}(y):=f(y-x)$. Concerning the generalization developed in this work, we consider an $n$-particle state $\Psi\in{{\cal H}}^{(n)}$ with $n>0$. For every Borel set $\Delta\subset\Sigma$, where $\Sigma$ is a spacelike 33-plane of Minkowski spacetime $\mathbb{M}$}, the effects $A_{f}^{u}(\Delta)$ of the notion of spatial localization we construct satisfy
@@
The family $A^{u}_{f}(\Delta)$ is a well-behaved positive-energy localization observable in each space ${{\cal H}}^{(n)}$, in particular for one-particle states. Notably, the integral for $\Delta=\Sigma$ is normalized independently of the choice of $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$, provided $\int_{\mathbb{M}}f^{2}d^{4}x=1$. The causality requirement denoted by $C$ and already established in [Mor23, DRM24] also holds for the localization observables constructed here. We further analyze the physical meaning of the notion of localization introduced above, showing that it is associated with the center of $u$-energy of the field and that, for one-particle states, it reduces to the observable already introduced in [Mor23]. In the one-particle case, one can say even more: if one takes $f(x^{0},\vec{x})=h^{\prime}(x^{0})h(\vec{x})$, then, in the non-relativistic large-mass limit, $h$ determines the precision of a von Neumann measurement scheme for detecting a single particle, whereas $h^{\prime}$ can be chosen independently.
@@ 1.3 Notions and notations
A. Minkowski spacetime. A four-dimensional real affine space whose space of translations $\mathsf{V}$ is equipped with a bilinear, non-degenerate, symmetric form $g$ with signature (,+,+,+)(-,+,+,+) is the Minkowski spacetime $\mathbb{M}$}. The points of $\mathbb{M}$} are called events and $g$ is called the Minkowski metric. We shall make use of the notation $u\cdot v:=g(u,v)$ if $u,v\in\mathsf{V}$. We also use the dot to indicate the standard (positive) scalar product of 33-vectors $\vec{u}\cdot\vec{v}$, viewing them as spacelike vectors (see below). Upon choosing an origin $o\in{\mathbb{M}}$, the points $p\in{\mathbb{M}}$ are in one-to-one correspondence with vectors of $\mathsf{V}$ through the map ${\mathbb{M}}\ni p\mapsto p-o\in\mathsf{V}$. We shall take advantage of this identification several times in the rest of this paper. If $p\in{\mathbb{M}}$ and $v\in\mathsf{V}$, $q=p+v$ means that qp=vq-p=v.
@@
If $A\subset{\mathbb{M}}$, the causal future $J^{+}(A):=\{p\in{\mathbb{M}}\>|\>p-q\in\overline{\mathsf{V}_{+}}\quad\mbox{for some $q\in A$}\}$ represents the events of $\mathbb{M}$} in the future of AA that can be physically influenced by AA. The causal past $J^{-}(A):=\{p\in{\mathbb{M}}\>|\>q-p\in\overline{\mathsf{V}_{+}}\quad\mbox{for some $q\in A$}\}$ is defined symmetrically.
@@
A rest space of a Minkowski reference frame $u\in{\mathsf{T}}_{+}$ is an affine 33-plane $\Sigma\subset{\mathbb{M}}$ $g$-normal to $u$, written $u\perp\Sigma$. If $o\in{\mathbb{M}}$ is a given origin, the family of rest spaces of $u\in{\mathsf{T}}_{+}$ is labeled by the time at which they occur in the said reference frame, $t_{o,\Sigma}=-(p-o)\cdot u$, which does not depend on $p\in\Sigma$. Every spacelike affine 33-plane is the rest space of a Minkowskian reference frame, indicated by $u_{\Sigma}$ and said to be adapted to $\Sigma$, at some time. $u_{\Sigma}$ is the future-directed unit normal vector to $\Sigma$, which is necessarily timelike.
@@
$\mathscr{B}(\Sigma)$ denotes the family of Borel subsets of a spacelike 33-plane $\Sigma$ and $\mathscr{B}_{b}(\Sigma)\subset\mathscr{B}(\Sigma)$ the subfamily of bounded elements, boundedness being equivalently referred to any of the said coordinate systems. $d\Sigma$ denotes the natural translationally invariant Borel measure on the rest space $\Sigma$ of $u_{\Sigma}$, which coincides with the Lebesgue measure on the ${\mathbb{R}}^{3}$ space of the spatial coordinates of any Minkowskian coordinate system $x^{0},x^{1},x^{2},x^{3}$ comoving with $u_{\Sigma}$. It is easy to prove that $d\Sigma$ does not depend on the choice of such a Minkowskian coordinate system. We use the notation $|\Delta|=\int_{\Sigma}\chi_{\Delta}(x)d\Sigma(x)=\int_{{\mathbb{R}}^{3}}\chi_{\Delta}(x)d^{3}x$ for the Lebesgue measure of $\Delta\in\mathscr{B}(\Sigma)$.
@@
Operators in Hilbert space have their own domains, $A:D(A)\to{{\cal H}}$, where $D(A)\subset{{\cal H}}$ is a subspace. The operations $A+B$, ABAB, aAaA ($a\in{\mathbb{C}}$) are defined on their standard domains: $D(A+B):=D(A)\cap D(B)$, $D(AB):=\{x\in D(B)\>|\>Bx\in D(A)\}$, $D(aA):=D(A)$ unless $a=0$, in which case $D(0A):=D(0):={{\cal H}}$. $A\geq 0$ means that $\langle\psi|A\psi\rangle\geq 0$ if $\psi\in D(A)$. In that case we say that the operator AA is positive.
@@ 2.1 Localization observables, causal conditions, non-commutativity
For future convenience, observe that, if $V\in{\mathfrak{B}}({{\cal H}})$ is a unitary operator and $({{\cal H}},{\cal R},A,U)$ is a relativistic spatial localization observable, then $({{\cal H}},{\cal R},A_{V},U_{V})$ is, where $A^{v}_{V}(\Delta):=VA^{v}(\Delta)V^{\dagger}$ and $(U_{V})_{g}:=VU_{g}V^{\dagger}$ is another relativistic spatial localization observable. It also satisfies the causality condition below and is of positive-energy type if $({{\cal H}},{\cal R},A,U)$ is.
@@
A relativistic spatial localization observable $({{\cal H}},{\cal R},A,U)$ is causal if it satisfies the following. For every $\Delta\in{\cal R}$, every spacelike 33-plane $\Sigma$, and every $v\in{\mathsf{S}}$,
@@
The explicit examples of relativistic spatial localization observables, especially of positive-energy type, presented in [Cas17, Mor23, DRM24, Cas24, Cas25, CDM26] for various types of particles satisfy $C$. They therefore show that Hegerfeldt’s causality issue can in fact be regarded as harmless when one considers certain unsharp notions of localization, whereas sharp ones are ruled out.
In [Mor26], after an accurate analysis of relativistic locality from Busch’s perspective, we have asserted that, dealing with relativistic spatial localization observable (also satisfying $C$), there is no compelling reason for requiring that $A^{v}(\Delta),A^{v^{\prime}}(\Delta^{\prime})]=0$ when $\Delta$ and $\Delta^{\prime}$ are sharply causally separated, i.e., they are included in respective open regions which are causally separated. This is true for systems like particles which have the propensity to localize in a unique position of a rest space when the rest space is ideally filled with detectors. The found result is in agreement with the celebrated achievement by Halvorson and Clifton [HaCl02] which proves that the above commutativity is actually forbidden if $({{\cal H}},{\cal R},A,U)$ is of positive-energy type [^4] This is even true with a weaker notion of localization observable than the one of Definition 2.2, see the review in [Mor26].. Failure of commutativity has an important consequence. If $\Delta$ and $\Delta^{\prime}$ are bounded regions contained in causally separated open sets $\cal O$} and $\cal O$$^{\prime}$ of $\mathbb{M}$}, one concludes that $A^{v}(\Delta)$ and $A^{v}(\Delta^{\prime})$ cannot belong to the corresponding local algebras of observables in the AHK approach. Indeed, if this were the case, they would necessarily commute. This shows that the effects $A^{v}(\Delta)$ of a relativistic spatial localization observable are not local observables in AHK sense.
@@ 2.2 Conditional localization and commutativity
If $\Sigma$ is a flat 33-dimensional plane, a laboratory $L(\Delta_{0})\subset{\mathbb{M}}$, with space given by a bounded set $\Delta_{0}\in\mathscr{B}(\Sigma)$, is defined as the causal completion $L(\Delta_{0}):=(\Delta_{0}^{\perp\!\!\!\!\perp})^{\perp\!\!\!\!\perp}$. If $\Delta_{0}$ is open in the relative topology, the resulting open set $L(\Delta_{0})$ is a globally hyperbolic spacetime in its own right, with $\Delta_{0}$ as a possible smooth spacelike Cauchy surface, as well as other curved smooth spacelike Cauchy surfaces.
@@
In other words, if we start with a state $\rho\in{\mathsf{S}}({{\cal H}})$ whose probability of finding the system in $\Delta_{0}$ is close to 11, $B^{V}_{\Delta_{0}}(\Delta)$ measures the fraction of detections in $\Delta$ over the total number of detections observed in the laboratory $\Delta_{0}$, when an ensemble of identical systems is prepared in the initial state $\rho$} and the relativistic spatial localization observable with effects ${\cal R}\ni\Delta\mapsto A_{V}(\Delta):=VA(\Delta)V^{\dagger}$ is used to define the spatial localization of the system. In this sense, $B^{V}_{\Delta_{0}}$ represents, at least in the above limit, a conditional localization observable: $B^{V}_{\Delta_{0}}(\Delta)$ gives the probability of finding the system in $\Delta\subset\Delta_{0}$, provided that it is known to be found in the bounded measurable set $\Delta_{0}$. This interpretation becomes increasingly accurate as the initial probability of finding the system in $\Delta_{0}$ approaches unity.
@@ 3.1 Bosonic Fock space and particle states
The latter identity is valid in every Minkowskian coordinate frame, where one identifies ${\mathbb{R}}^{3}$ with the 33-space of the spatial part of the four-momenta, and